You receive a prescription written for the patient to take 2 tablets Q 8 hours, dispense 180 tablets. How long will this prescription last? Calculate the days supply.
1
Expert's answer
2012-11-09T06:13:01-0500
2*24/8=3*2=6 tablets a day
180/6=30 days.
So, prescription will last for30 days